Okay THe question is there is a 50 foot flag pole that has snapped and fallen down and the tip lays 40 feet from the base, how much of the pole is left standing * the pictures shows a right triangle,, i have no clue how to solve this problem..
Answer by ankor@dixie-net.com(22740)

there is a 50 foot flag pole that has snapped and fallen down and the tip lays 40 feet from the base, how much of the pole is left standing
:
This is a right triangle problem a^2 + b^2 = c^2
Let a = amt of pole left standing
Let b = 40 ft distance from the base of the pole to the tip of the fallen part
Let c = the hypotenuse (the broken part from the pole to the ground
:
a^2 + 40^2 = c^2
:
The pole is 50' high, therefore
c = (50-a)
therefore
a^2 + 40^2 = (50-a)^2
a^2 + 1600 = 2500 - 100a + a^2
a^2 - a^2 + 100a = 2500 - 1600
100a = 900
a = 900/100
a = 9 ft of the pole is left standing
:
:
Check this, the hypotenuse will be 50 - 9 = 41 ft
9^2 + 40^2 = 41^2
81 + 1600 = 1681; confirms our solution of a = 9